# Multiple dispersive bounds. II) Sub-threshold branch-cuts

### v2

@@ Abstract
We apply the general framework of multiple dispersive bounds, firstly discussed in the companion paper [1], to the study of sub-threshold branch-cuts. We propose the simultaneous application of a double dispersive bound as a proper way to take into account unitarity constraints within phenomenological analyses of hadronic form factors in the presence of sub-threshold branch-cuts. Accordingly, the standard $z$-expansion of hadronic form factors, commonly referred to as the Boyd-Grinstein-Lebed approach [2, 3, 4, 5], is modified by including simultaneously the dispersive bounds related to the pair-production and to the sub-threshold regions. For the latter one the effects of above-threshold poles are described through a new resonance model and the possible choices of the outer function outside the pair-production region are discussed. A detailed numerical analysis of the experimental data or lattice QCD results in the spacelike region for the charged kaon form factor is presented as a direct application of the procedure of double dispersive bound. The comparison with other methodologies present in literature and with the $z$-expansion based on the single, total dispersive bound clearly shows that the $z$-expansion including the double dispersive bound provides the most precise extrapolation at large momentum transfer as well as the most stable results with respect to the choice of the outer function outside the pair-production region.
@@ I Introduction
In this work, we apply the general framework of multiple dispersive bounds, firstly discussed in the companion paper [1], to the study of sub-threshold branch-cuts within the standard $z$-expansion of hadronic form factors, commonly referred to as the Boyd-Grinstein-Lebed (BGL) approach [2, 3, 4, 5]. Our goal is to implement the properties of unitarity and analyticity by imposing a double dispersive bound, which takes simultaneously into account unitarity in the pair-production region and within the lowest sub-threshold and the pair-production branch-points. We stress the importance of considering both constraints in the BGL $z$-expansion of hadronic form factors, at variance with what has been done till now in literature. In particular, while the first bound can be imposed through the knowledge of the so-called susceptibility, which can be estimated from the derivatives of appropriate two-point Green functions evaluated in momentum space, the second bound requires a separate estimate in a range of timelike values of the momentum transfer not accessible by experiments. To this end we develop a model for the effects of above-threshold resonances on the hadronic form factor. Our model successfully describes the $\rho$(770)-meson resonance in the case of the electromagnetic form factor of the pion, as well as the main poles entering the study of the electromagnetic form factors of charged and neutral kaons. Furthermore, we highlight that the choice of the outer function outside the pair-production arc is not unique and possible choices are discussed.
@@ II.1 A model for the extra branch-cut
Let us indicate with $f^{th}(z_{+})$ the contribution of the single extra branch-cut, starting at $t_{th}<t_{+}$, to the form factor $f(z_{+})$. For real values of $z_{+}$ above $z_{+}^{th}$ the function $f^{th}(z_{+})$ is real, while for $z_{+}$ below $z_{+}^{th}$ it acquires an imaginary part, which represents (half of) the discontinuity of the form factor $f(z_{+})$ along the extra branch-cut. Therefore, the difference $f(z_{+})-f^{th}(z_{+})$ is an analytic function of the real type inside the unit disk $|z_{+}|<1$, where it has a Taylor expansion with real coefficients. One gets
@@ II.3 Semi-inclusive dispersive bounds
The use of a finite order $M$ in the truncated expansion (31) arises the issue of truncation errors, which were firstly discussed in Refs. [4, 5]. There, an upper limit to the truncation error was obtained, assuming that the coefficients $a_{k}^{(M)}$ coincide with the first $(M+1)$ coefficients $a_{k}$ of the untruncated BGL expansion (12). See for details Appendix B.
@@ II.4 Differences with other expansions
In terms of the monomials $z^{k}$ the $z$-expansion (35) can be written as (see Appendix C)
@@ III.1 Effects of above-threshold poles on the form factor $f(t)$
where $f_{R}(0)\equiv f_{R}(z_{+}=0)$ represents the value of the form factor at $z_{+}=0$ (i.e. at $t=t_{0}$) and the term $(1-z_{+})^{2}$ in the numerator has been inserted in order to guarantee that $f_{R}(z)\propto 1/(-t)$ at large values of $|t|$. In Eq. (43) the conjugate poles occur only in the second Riemann sheet ($|z_{+}|>1$) at $z_{+}=1/z_{R}^{\pm}$. The function $f_{R}(z_{+})$ is analytic of the real type inside the unit disk $|z_{+}|<1$ and requires only the knowledge of the mass $M_{R}$ and width $\Gamma_{R}$ of the resonance.
@@ III.2 The electromagnetic form factor of the charged pion
We now check the quality of the approximation (43) by considering the case of the electromagnetic form factor for charged pions, $F_{\pi}^{(em)}(t)$, which is known to receive a dominant contribution from the $\rho(770)$-meson resonance for $|t|\lesssim 1$ GeV2 (see also the recent work of Ref. [34]). In this note we make use of the results of Ref. [35], based on a detailed analysis of a large set of both space-like ($t<0$) and time-like experimental data up to $t=1$ GeV2, carried out within a dispersive approach satisfying unitarity and analyticity.
Since the pion form factor $F_{\pi}^{(em)}(t)$ is related to the matrix element of a vector current, angular momentum conservation imposes a specific constraint at the threshold $t_{+}=4M_{\pi}^{2}$, namely, in terms of the conformal variable $z_{+}$ one must have $dF_{\pi}^{(em)}(z_{+})/dz_{+}\Big|_{z+=-1}=0$ (see Ref. [36]). Therefore, we modify Eq. (43) with $R=\rho$ by including as extra factor $(1+b_{1}z_{+})$, with $b_{1}$ chosen to guarantee that $df_{\rho}(z_{+})/dz_{+}\Big|_{z_{+}=-1}=0$. Moreover, charge conservation implies that $F_{\pi}^{(em)}(t=0)=1$, which can be easily implemented by choosing the auxiliary variable $t_{0}$ in the definition of the conformal variable $z_{+}$ equal to $t_{0}=0$ and by putting $f_{\rho}(z_{+}=0)=1$. One has
@@ III.3 A modified $z$-expansion to include effects from above-threshold poles
Since $f_{R}(z_{+})$ is an analytic function of the real type inside the unit disk $|z_{+}|<1$, an interesting development could be to modify the standard BGL $z$-expansion (12) or the Bourrely-Caprini-Lellouch (BCL) $z$-expansion [36] by taking into account an above-threshold resonance through the function (43).
@@ IV The electromagnetic form factor of the charged kaon
We now basically refer to the dispersive analysis made in Ref. [43]. The electromagnetic form factors of charged and neutral kaons, $F_{K\pm}^{(em)}(t)$ and $F_{K^{0}}^{(em)}(t)$, can be decomposed into an isovector (V) and an isoscalar (S) contributions
@@
Data on the phase and modulus of $g_{1}^{1}(t)$ are available in the physical region $t\geq t_{2K}$ from $\pi\pi\to\overline{K}K$ scattering. Instead, in the unphysical region $t_{2\pi}\leq t\leq t_{2K}$ such data cannot occur. However, thanks to Eq. (58) the phase of $g_{1}^{1}(t)$ must compensate the one of the pion form factor, which in turn for $t\leq t_{2K}\sim 1$ GeV2 is dominated by the phase of the elastic $\pi\pi$ scattering. Since the phase of $g_{1}^{1}(t)$ is known for $t\geq t_{2\pi}$, then its modulus can be reconstructed [45] using the standard Muskhelishvili-Omnès method [46, 42]. Using dispersive parameterizations of $g_{1}^{1}(t)$, elaborated in Ref. [45], and of the pion form factor $F_{\pi}^{(em)}(t)$ (similar to the one developed in Ref. [35]), the isovector part of the kaon form factor $F_{K}^{(V)}(t)$ was determined in Ref. [43] for $t\geq t_{2\pi}$ and tested successfully against experimental data on $\tau^{-}\to K^{-}K_{S}\nu_{\tau}$ decays [47] for $t\geq t_{2K}$. It turned out that $F_{K}^{(V)}(t)$ receives an important contribution from the $\rho(770)$-meson resonance and to a lesser extent from the $\rho^{\prime}(1450)$-meson resonance.
@@ V The outer function $\phi(z)$ outside the pair-production arc
where the function $\phi_{+}(z_{+})$ is the outer function adopted in the BGL approach [2, 3, 4, 5] to describe the momentum dependence of the form factors relevant for the semileptonic weak decays of hadrons [5, 48] (see Appendix A of the companion paper [1] for the case of semileptonic decays of pseudoscalar mesons). In this way, for obtaining an upper limit to the contribution $\chi_{+}^{U}$ inside the pair-production arc one can make use of the susceptibility $\chi_{n}$, related to the appropriate derivative of the vacuum polarization function and calculable non-perturbatively on the lattice (see Refs. [49, 50, 51, 18, 52]). We stress that this can be done without any knowledge of the $z$-dependence of the form factor on the pair-production arc.
@@ VI Analysis of the spacelike experimental and lattice data on the charged kaon form factor
We want to compare the results obtained using the truncated BGL expansion (31) applying either the double bound represented by Eqs. (32)-(33) or the single, total bound (34). In addition we apply also the expansion (37) with the unitarity constraint (38), as described in Section II.4. We remind that such an expansion, adopted in Refs. [24, 25, 26], is equivalent to an expansion of the form factor in terms of Szegő polynomials, proposed in Refs. [21, 22, 23], but it implements the unitarity constraint in an incomplete way, i.e., only on the pair-production arc.
@@
In the case of the experimental data on $F_{K^{\pm}}^{(em)}(Q^{2})$ there are 10 data from FNAL [6] and 15 data from CERN [7]. To ensure charge conservation we add a further data point $F_{K^{\pm}}^{(em)}(Q^{2}=0)=1$ with a tiny error ($\simeq 10^{-12}$) [^13] Choosing $t_{0}=t_{-}=0$ in Eq. (11), this is equivalent to fix the coefficients $a_{0}$ of the BGL expansion (31) to the value $a_{0}=\phi(z=0)/\sqrt{\chi^{U}}$.. Therefore, we have a total of $N_{data}=26$ data points. The covariance matrices of the two experiments are known separately, so that we can construct the full covariance matrix $C_{ij}$ of the input data by considering the two experiments uncorrelated.
@@
Let us start with the case of the experimental data on the charged kaon form factor and the choice $q=0$ for the kinematical function outside the pair-production arc. We apply the truncated BGL expansion (31) at various order $M$ of the truncation using either the double bound (32)-(33) or the single, total bound (34). The dispersive bounds $\chi_{+}^{U}$ and $\chi_{\rm extra}^{U}$ are given by Eqs. (71)-(72) and, therefore, at $q=0$ the total bound $\chi^{U}=\chi_{+}^{U}+\chi_{\rm extra}^{U}$ is largely dominated by the pair-production bound $\chi_{+}^{U}$. We observe a nice stability of the extrapolated form factor up to $Q^{2}\simeq 1$ GeV2 for $M\geq 6$. We apply also the expansion (37) with the unitarity constraint (38), restricted to the pair-production arc only. A reasonable stability of the results for the form factor up to $Q^{2}\simeq 1$ GeV2 is observed against the order $M$ of the truncation for $M\geq 6$.
The comparison of the predicted bands for the form factor up to $Q^{2}\simeq 1$ GeV2 is shown in Fig. 5, where we have truncated all the $z$-expansions at order $M=8$. In what follows the width of the bands will be always provided at $1\sigma$ level. For the three expansions the value of the reduced (correlated) $\chi^{2}$-variable turns out to be equal to $\chi^{2}/(d.o.f.)\simeq 1.3$.
@@
  • The use of the expansion (37) with the pair-production bound (38) alone, leads to a quite huge spread of values at large $Q^{2}$. This is related to the fact that the expansion in terms of Szegő polynomials, proposed in Refs. [21, 22, 23] and adopted also in Refs. [24, 25, 26], represents an incomplete application of the unitarity constraints and it may produce violations of the unitarity constraint outside the pair-production arc and, consequently, of the total bound $\chi^{U}$ on the full arc. Indeed, in the case at hand we observe that the coefficients $b_{k}$ of the expansion (37) have large numerical values, leading to $\sum_{k=0}^{M}b_{k}^{2}>>\chi^{U}/\chi_{+}^{U}\simeq 1.1$.
@@
We observe again that:   i) the use of the double bound improves the extrapolation at large $Q^{2}$ by a factor of $\sim 2$ with respect to the use of the single, total bound;   ii) the two BGL expansions implementing properly the unitarity constraints are nicely consistent with the lattice QCD data up to $Q^{2}\simeq 28$ GeV2, and   iii) the expansion in terms of Szegő polynomials, which takes into account only the unitarity constraint on the pair-production arc, is only marginally consistent with the lattice QCD data at large $Q^{2}$.
We now move to the other choice considered for the kinematical function $\phi^{q_{1}q_{2}}(z)$ , namely $q=q_{1}+q_{2}=3.5$. For the bound $\chi_{\rm extra}^{U}$ we use the estimate given by Eq. (73), which is much larger than the pair-production contribution $\chi_{+}^{U}$, given by Eq. (71), by a factor of $\approx 240$. Thus, at $q=3.5$ the total bound $\chi^{U}=\chi_{+}^{U}+\chi_{\rm extra}^{U}$ is largely dominated by the contribution $\chi_{\rm extra}^{U}$ outside the pair-production arc. In what follows we will limit ourselves to the two BGL expansions properly compatible with unitarity constraints, i.e. those based on Eq. (31) with either the double bound (32)-(33) or the single, total bound (34).
@@
Both BGL expansions, including only lattice points with $Q^{2}\lesssim 0.4$ GeV2, predict values consistent with the lattice QCD data up to $Q^{2}\simeq 28$ GeV2 though within a much larger band of values with respect to the choice $q=0$ (compare with Fig. 6). Moreover, an anomalous increase of the width of the predicted bands is visible also at low $Q^{2}$, i.e. in the $Q^{2}$-range of values adopted in the fitting procedure. Both unpleasant features are related to the presence of non-unitary effects in the set of input data. The removal of such effects is discussed in the next subsection.
VI.1 Unitary filtering of the input data

### v3

@@ Abstract
We apply the strategy proposed in the companion paper [1] for dealing with multiple dispersive bounds, to the case of sub-threshold branch-cuts, which is a topic addressed extensively in the literature (see, e.g., Refs. [2, 3, 4, 5, 6, 7, 8]). We consider the simultaneous application of a double dispersive bound as a proper way to take into account unitarity constraints within phenomenological analyses of hadronic form factors in the presence of sub-threshold branch-cuts. Accordingly, the standard $z$-expansion of hadronic form factors, commonly referred to as the Boyd-Grinstein-Lebed approach [9, 10, 4, 11], is modified by including simultaneously the dispersive bounds related to the pair-production and to the sub-threshold regions. For the latter one the effects of above-threshold poles are described through a simple resonance model and the possible choices of the outer function outside the pair-production region are discussed. A detailed numerical analysis of the experimental data or lattice QCD results in the spacelike region for the charged kaon form factor is presented as a direct application of the procedure of double dispersive bound. The comparison with other methodologies present in literature and with the $z$-expansion based on the single, total dispersive bound clearly shows that the $z$-expansion including the double dispersive bound provides the most precise extrapolation at large momentum transfer as well as the most stable results with respect to the choice of the outer function outside the pair-production region.
@@ I Introduction
The topic of sub-threshold branch-cuts has been addressed extensively in the literature (see, e.g., Refs. [2, 3, 4, 5, 6, 7, 8]). In this work, we apply the strategy proposed in the companion paper [1] for dealing with multiple dispersive bounds, to the case of of sub-threshold branch-cuts within the standard $z$-expansion of hadronic form factors, commonly referred to as the Boyd-Grinstein-Lebed (BGL) approach [9, 10, 4, 11]. Our goal is to implement the properties of unitarity and analyticity by imposing a double dispersive bound, which takes simultaneously into account unitarity in the pair-production region and within the lowest sub-threshold and the pair-production branch-points. We stress the importance of considering both constraints in the BGL $z$-expansion of hadronic form factors, as already attempted in Refs. [3, 4, 12, 8], but at variance with what has been done recently in literature (see later Section II.4). In particular, while the first bound can be imposed through the knowledge of the so-called susceptibility, which can be estimated from the derivatives of appropriate two-point Green functions evaluated in momentum space, the second bound requires a separate estimate in a range of timelike values of the momentum transfer not accessible by experiments. To this end we develop a model for the effects of above-threshold resonances on the hadronic form factor. Our model successfully describes the $\rho$(770)-meson resonance in the case of the electromagnetic form factor of the pion, as well as the main poles entering the study of the electromagnetic form factors of charged and neutral kaons. Furthermore, we highlight that the choice of the outer function outside the pair-production arc is not unique and possible choices are discussed.
@@ II.1 A model for the extra branch-cut
Let us indicate with $f^{th}(z_{+})$ the contribution of the single extra branch-cut, starting at $t_{th}<t_{+}$, to the form factor $f(z_{+})$. For real values of $z_{+}$ above $z_{+}^{th}$ the function $f^{th}(z_{+})$ is real, while for $z_{+}$ below $z_{+}^{th}$ it acquires an imaginary part, which represents (half of) the discontinuity of the form factor $f(z_{+})$ along the extra branch-cut. Therefore, the difference $f(z_{+})-f^{th}(z_{+})$ is an analytic function of the real type inside the unit disk $|z_{+}|<1$, where it has a Taylor expansion with real coefficients. One gets [2, 3, 4]
@@ II.3 Semi-inclusive dispersive bounds
The use of a finite order $M$ in the truncated expansion (31) raises the issue of truncation errors, which were firstly discussed in Refs. [4, 11]. There, an upper limit to the truncation error was obtained, assuming that the coefficients $a_{k}^{(M)}$ coincide with the first $(M+1)$ coefficients $a_{k}$ of the untruncated BGL expansion (12). See for details Appendix B.
@@ II.4 Differences with other expansions
As already shown in Appendix B of Ref. [29] in terms of the monomials $z^{k}$ the $z$-expansion (35) can be written as (see also Appendix C)
@@ III.1 Effects of above-threshold poles on the form factor $f(t)$
where $f_{R}(0)\equiv f_{R}(z_{+}=0)$ represents the value of the form factor at $z_{+}=0$ (i.e. at $t=t_{0}$) and the term $(1-z_{+})^{2}$ in the numerator has been inserted in order to guarantee that $f_{R}(z)\propto 1/(-t)$ at large values of $|t|$. We stress that Eq. (43) is similar to the ansatz discussed in Ref. [36]: the resonance $R$ produces conjugate poles occurring only in the second Riemann sheet ($|z_{+}|>1$) at $z_{+}=1/z_{R}^{\pm}$, and the function $f_{R}(z_{+})$ is analytic of the real type inside the unit disk $|z_{+}|<1$. Note that in this way only the knowledge of the mass $M_{R}$ and width $\Gamma_{R}$ of the resonance is required.
@@ III.2 The electromagnetic form factor of the charged pion
We now check the quality of the approximation (43) by considering the case of the electromagnetic form factor for charged pions, $F_{\pi}^{(em)}(t)$, which is known to receive a dominant contribution from the $\rho(770)$-meson resonance for $|t|\lesssim 1$ GeV2. In this work we make use of the results of Ref. [39], based on a detailed analysis of a large set of both space-like ($t<0$) and time-like experimental data up to $t=1$ GeV2, carried out within a dispersive approach satisfying unitarity and analyticity.
Since the pion form factor $F_{\pi}^{(em)}(t)$ is related to the matrix element of a vector current, angular momentum conservation imposes a specific constraint at the threshold $t_{+}=4M_{\pi}^{2}$, namely, in terms of the conformal variable $z_{+}$ one must have $dF_{\pi}^{(em)}(z_{+})/dz_{+}\Big|_{z+=-1}=0$ (see Ref. [40]). Therefore, we modify Eq. (43) with $R=\rho$ by including as extra factor $(1+b_{1}z_{+})$, with $b_{1}$ chosen to guarantee that $df_{\rho}(z_{+})/dz_{+}\Big|_{z_{+}=-1}=0$. Moreover, charge conservation implies that $F_{\pi}^{(em)}(t=0)=1$, which can be easily implemented by choosing the auxiliary variable $t_{0}$ in the definition of the conformal variable $z_{+}$ equal to $t_{0}=0$ and by putting $f_{\rho}(z_{+}=0)=1$. One has (see also the similar ansatz adopted in Ref. [41])
@@ III.3 A modified $z$-expansion to include effects from above-threshold poles
Since $f_{R}(z_{+})$ is an analytic function of the real type inside the unit disk $|z_{+}|<1$, an interesting development, considered already in Ref. [36], is to modify the standard BGL $z$-expansion (12) or the Bourrely-Caprini-Lellouch (BCL) $z$-expansion [40] by taking into account an above-threshold resonance through the function (43).
@@ IV The electromagnetic form factor of the charged kaon
We now basically refer to the dispersive analysis made in Ref. [48]. The electromagnetic form factors of charged and neutral kaons, $F_{K\pm}^{(em)}(t)$ and $F_{K^{0}}^{(em)}(t)$, can be decomposed into isovector (V) and an isoscalar (S) contributions
@@
Experimental data on the phase and modulus of $g_{1}^{1}(t)$ are available in the physical region $t\geq t_{2K}$ from $\pi\pi\to\overline{K}K$ scattering. Instead, in the unphysical region $t_{2\pi}\leq t\leq t_{2K}$ such data are not experimentally accessible. However, thanks to Eq. (58) the phase of $g_{1}^{1}(t)$ must compensate the one of the pion form factor, which in turn for $t\leq t_{2K}\sim 1$ GeV2 is dominated by the phase of the elastic $\pi\pi$ scattering. Since the phase of $g_{1}^{1}(t)$ is known for $t\geq t_{2\pi}$, then its modulus can be reconstructed [50] using the standard Muskhelishvili-Omnès method [51, 47]. Using dispersive parameterizations of $g_{1}^{1}(t)$, elaborated in Ref. [50], and of the pion form factor $F_{\pi}^{(em)}(t)$ (similar to the one developed in Ref. [39]), the isovector part of the kaon form factor $F_{K}^{(V)}(t)$ was determined in Ref. [48] for $t\geq t_{2\pi}$ and tested successfully against experimental data on $\tau^{-}\to K^{-}K_{S}\nu_{\tau}$ decays [52] for $t\geq t_{2K}$. It turned out that $F_{K}^{(V)}(t)$ receives an important contribution from the $\rho(770)$-meson resonance and to a lesser extent from the $\rho^{\prime}(1450)$-meson resonance.
@@ V The outer function $\phi(z)$ outside the pair-production arc
where the function $\phi_{+}(z_{+})$ is the outer function adopted in the BGL approach [9, 10, 4, 11] to describe the momentum dependence of the form factors relevant for the semileptonic weak decays of hadrons [11, 53] (see Appendix A of the companion paper [1] for the case of semileptonic decays of pseudoscalar mesons). In this way, for obtaining an upper limit to the contribution $\chi_{+}^{U}$ inside the pair-production arc one can make use of the susceptibility $\chi_{n}$, related to the appropriate derivative of the vacuum polarization function and calculable nonperturbatively on the lattice (see Refs. [54, 55, 56, 24, 57]). We stress that this can be done without any knowledge of the $z$-dependence of the form factor on the pair-production arc.
@@ VI Analysis of the spacelike experimental and lattice data on the charged kaon form factor
We want to compare the results obtained using the truncated BGL expansion (31) applying either the double bound represented by Eqs. (32)-(33) or the single, total bound (34). In addition we apply also the expansion (37) with the unitarity constraint (38), as described in Section II.4. We remind that such an expansion, adopted in Refs. [29, 30, 31], is equivalent to the expansion of the form factor in terms of Szegő polynomials, proposed in Refs. [26, 27, 28], namely Eq. (35) with the constraint (36). This approach implements unitarity in an incomplete way, i.e., only on the pair-production arc, and it does not guarantee unitarity on the full circle.
@@
In the case of the experimental data on $F_{K^{\pm}}^{(em)}(Q^{2})$ there are 10 data from FNAL [13] and 15 data from CERN [14]. To ensure charge conservation in the BGL $z$-expansion (31) we add a further data point $F_{K^{\pm}}^{(em)}(Q^{2}=0)=1$ with a tiny error ($\simeq 10^{-12}$) [^13] Choosing $t_{0}=t_{-}=0$ in Eq. (11), this is equivalent to fix the coefficient $a_{0}$ of the BGL expansion (31) to the value $a_{0}=\phi(z=0)/\sqrt{\chi^{U}}$ without introducing the additional point at $Q^{2}=0$.. Therefore, we have a total of $N_{data}=26$ data points. The covariance matrices of the two experiments are known separately, so that we can construct the full covariance matrix $C_{ij}$ of the input data by considering the two experiments uncorrelated.
@@
In Section VI.1 we perform the BGL analyses without applying any unitary filtering to the input datasets, namely the BGL $z$-expansions are applied to samples of input data containing both non-unitary and unitary events. Instead, in Section VI.2 we filter the input datasets against the unitarity bound(s) before applying the BGL $z$-expansions.
VI.1 Unfiltered input data
Let us start with the case of the experimental data on the charged kaon form factor and the choice $q=0$ for the kinematical function outside the pair-production arc. We apply the truncated BGL expansion (31) at various order $M$ of the truncation using either the double bound (32)-(33) or the single, total bound (34). The dispersive bounds $\chi_{+}^{U}$ and $\chi_{\rm extra}^{U}$ are given by Eqs. (71)-(72) and, therefore, at $q=0$ the total bound $\chi^{U}=\chi_{+}^{U}+\chi_{\rm extra}^{U}$ is largely dominated by the pair-production bound $\chi_{+}^{U}$. We observe a nice stability of the extrapolated form factor up to $Q^{2}\simeq 1$ GeV2 for $M\geq 6$. We apply also the expansion (37) with the unitarity constraint (38), restricted to the pair-production arc only.
The comparison of the predicted bands for the form factor up to $Q^{2}\simeq 1$ GeV2 is shown in Fig. 5. In what follows the width of the bands will be always provided at $1\sigma$ level. We have checked that:   i) the probability distributions of the calculated form factors are largely normal, and   ii) the calculated bands (i.e, both central values and errors) are stable against variation of the truncation order $M$. We have found that the stability starts already at $M=4$. In Fig. 5 all the $z$-expansions are truncated at order $M=8$. For the three expansions the value of the reduced (correlated) $\chi^{2}$-variable turns out to be equal to $\chi^{2}/(d.o.f.)\simeq 1.3$.
@@ VI.1 Unfiltered input data
  • The use of the expansion (37) with the pair-production bound (38) alone, leads to a quite huge spread of values at large $Q^{2}$. This is related to the fact that the expansion in terms of Szegő polynomials, proposed in Refs. [26, 27, 28] and adopted also in Refs. [29, 30, 31], represents an incomplete application of the unitarity constraints and it may produce violations of the unitarity constraint outside the pair-production arc and, consequently, of the total bound $\chi^{U}$ on the full arc. Indeed, in the case at hand we observe that the coefficients $b_{k}$ of the expansion (37) have large absolute values, leading to $\sum_{k=0}^{M}b_{k}^{2}>>\chi^{U}/\chi_{+}^{U}\simeq 1.1$.
@@
We observe again that:   i) the use of the double bound improves the extrapolation at large $Q^{2}$ by a factor of $\sim 2$ with respect to the use of the single, total bound;   ii) the two BGL expansions implementing properly the unitarity constraints are consistent with the lattice QCD data up to $Q^{2}\simeq 28$ GeV2, and   iii) the expansion in terms of Szegő polynomials, which takes into account only the unitarity constraint on the pair-production arc, is consistent with the lattice QCD data at large $Q^{2}$, i.e. in the region where the LQCD data points are not used in the fitting procedure; however, the coefficients $b_{k}$ of the expansion (37) turn out to have large absolute values, leading to $\sum_{k=0}^{M}b_{k}^{2}>>\chi^{U}/\chi_{+}^{U}\simeq 1.1$, which implies that unitarity on the full arc is badly violated.
We now move to the other choice considered for the kinematical function $\phi^{q_{1}q_{2}}(z)$ , namely $q=q_{1}+q_{2}=3.5$. For the bound $\chi_{\rm extra}^{U}$ we use the estimate given by Eq. (73), which is much larger than the pair-production contribution $\chi_{+}^{U}$, given by Eq. (71), by a factor of $\approx 240$. Thus, at $q=3.5$ the total bound $\chi^{U}=\chi_{+}^{U}+\chi_{\rm extra}^{U}$ is largely dominated by the contribution $\chi_{\rm extra}^{U}$ outside the pair-production arc. In what follows, we will leave out the results corresponding to the expansion (35), expressed in terms of Szegő polynomials. As previously shown at $q=0$ and checked also at $q=3.5$, this approach violates badly unitarity on the full arc, introducing therefore non-unitarity effects already at the level of the $z$-expansion. Our aim is to focus on the results of the two BGL $z$-expansions properly compatible with unitarity, i.e. those based on Eq. (31) with either the double bound (32)-(33) or the single, total bound (34).
@@
Both BGL expansions, including only lattice points with $Q^{2}\lesssim 0.4$ GeV2, predict values consistent with the lattice QCD data up to $Q^{2}\simeq 28$ GeV2 though within a much larger band of values with respect to the choice $q=0$ (compare with Fig. 6). Moreover, at variance with Fig. 6, a large width of the predicted bands is visible also at low $Q^{2}$, i.e. in the $Q^{2}$-range of values adopted in the fitting procedure. Both unpleasant features are related to the presence of non-unitary effects in the set of input data. The removal of such effects is discussed in the next subsection.
VI.2 Unitary filtering of the input data